Could you please show me how to work this problem so I know how to do it.
Perimeter of a rectangle. The perimeter of a rectangular backyard is 6x + 6 yards. If the width is x yards, find the binomial that represents the length.
Thank you

Answer by mananth(16949)

The perimeter of a rectangular backyard is 6x + 6 yards. If the width is x yards, find the binomial that represents the length.
...
Perimeter = 6x+6 yards
width = x
..
2*(L+W)=perimeter
2*(L+x)= 6x+6
/2
L+x= 3x+3
-x
L+x-x=3x+3-x
L= 2x+3 yards
...
m.ananth@hotmail.ca
...
CHECK
2(2x+3)+2x = Perimeter
4x+6+2x
6x+6